# Changzhi Sun, Jixiang Luo, Dell Zhang#, Hao Sun, Chi Zhang, Xuelong Li#

### v2

@@ Abstract
Egocentric AI assistants in real-world settings must process multi-modal inputs (video, audio, text), respond in real time, and retain evolving long-term memory. However, existing benchmarks typically evaluate these abilities in isolation, lack realistic streaming scenarios, or support only short-term tasks. We introduce TeleEgo, a long-duration, streaming, omni-modal benchmark for evaluating egocentric AI assistants in realistic daily contexts. The dataset features over 14 hours per participant of synchronized egocentric video, audio, and text across four domains: work & study, lifestyle & routines, social activities, and outings & culture. All data is aligned on a unified global timeline and includes high-quality visual narrations and speech transcripts, curated through human refinement. TeleEgo defines 12 diagnostic subtasks across three core capabilities: Memory (recalling past events), Understanding (interpreting the current moment), and Cross-Memory Reasoning (linking distant events). It contains 3,291 human-verified QA items spanning multiple question formats (single-choice, binary, multi-choice, and open-ended), evaluated strictly in a streaming setting. We propose two key metrics — Real-Time Accuracy and Memory Persistence Time — to jointly assess correctness, temporal responsiveness, and long-term retention. TeleEgo provides a realistic and comprehensive evaluation to advance the development of practical AI assistants.
@@ 1 Introduction
To address these challenges, we present TeleEgo: a long-duration, streaming, and fully multimodal benchmark grounded in real-world scenarios, purpose-built to evaluate egocentric AI assistants (see Figure 1). TeleEgo consists of synchronized video, audio, and textual data collected from multiple participants, each contributing more than 14 hours of recordings. The dataset spans four major domains: work & study, lifestyle & routines, social activities, and outings & culture. All data streams are precisely aligned to a unified global timeline and enriched with manually curated speech transcripts and visual narrations to ensure high quality and semantic clarity. Building on this foundation, TeleEgo introduces 12 diagnostic tasks covering three core ability: memory (recalling past events), understanding (interpreting the present), and cross-memory reasoning (connecting distant moments). These tasks include a total of 3,291 human-verified QA items across various formats, including single-choice, binary, multiple-choice, and open-ended questions. Each task is tied to a specific time point and a decision window, requiring models not only to answer correctly, but also to respond at the right time. All evaluations are conducted under streaming conditions, and a task is considered successful only if the model’s first correct answer falls within the allowed window. This setup prevents models from guessing too early or answering too late, offering a more accurate measure of real-time decision-making and responsiveness. We further introduce two key evaluation metrics: Real-Time Accuracy, which captures the model’s ability to produce correct responses within the given time constraint, and Memory Persistence Time, which measures how long a model can retain past information without re-encountering the original context. These metrics offer a comprehensive view of not just correctness, but also temporal responsiveness and long-term memory retention.
@@
  • An evaluation protocol tailored to streaming inputs, using two core metrics, real-time accuracy and memory persistence, to assess comprehensive performance.
2 Related Work
Egocentric Models and Benchmarks. Egocentric (first-person) vision research has grown significantly, evolving from early single-user recordings to diverse, large-scale datasets capturing daily life from a first-person perspective. Pioneering work such as EPIC-KITCHENS [5] and its extensions like VISOR [6] introduced large-scale, object-rich video datasets in home environments. Later, Ego4D and Ego-Exo4D [8, 9] broadened the scope to include tasks like episodic memory, future prediction, and skill learning, incorporating both egocentric and exocentric views. Recent datasets have begun to explore more complex cognitive tasks and assistant-like interactions. EgoLife [21] captures long-duration recordings from a single household, while EgoThink [3] and EgoExoLearn [10] focus on segment-level reasoning and teaching-following dynamics. Other work like MM-Ego [22] and EgoTextVQA [25] targets memory and text-based understanding. However, most of these evaluations remain offline and task-specific. TeleEgo advances egocentric benchmarking by combining multi-day, multi-role, and multi-theme recordings with dual text annotations aligned to a unified global timeline. Its online evaluation protocol measures not only real-time decision-making, but also the persistence of memory over time—enabling more realistic and comprehensive evaluations of assistant capabilities.
@@ 2 Related Work
Omni-Modal Assistants. Recent advancements in omni-modal models aim to unify understanding across multiple input types—such as text, vision, audio, and speech—enabling more flexible and human-like assistants. Closed-source systems like GPT-4o [12] and Gemini 1.5/2.5 [18, 4] support end-to-end speech and video processing with long multimodal context windows, allowing for rich, coherent interactions. Open-source progress has also accelerated. Models such as LLaVA–OneVision [13] and InternVL-2.5 [2] support multi-image and video understanding. Others like Qwen2.5–Omni [20], MiniCPM–o [19], VITA [7], and Baichuan–Omni [14] are optimized for real-time, streaming audiovisual input and response. Despite their strong performance, these models are typically evaluated in offline, task-specific settings. Few benchmarks assess whether assistants can respond accurately within time-sensitive decision windows or retain relevant information across long temporal spans. By combining egocentric recordings, streaming video contexts, and real-time memory evaluation, TeleEgo provides a unified testbed to study omni-modal assistant performance in realistic environments—bridging the gap between research in egocentric datasets, streaming benchmarks, and multimodal models.
@@ 3.4 Benchmark QA Annotation
As shown in Figure 4, the TeleEgo benchmark is hierarchically structured along three key cognitive dimensions—memory, understanding, and reasoning—with twelve subcategories providing finer granularity. The final dataset contains 3,291 verified QA instances across all four formats (see Table 2). Subcategory distributions demonstrate balanced coverage across temporal, causal, and semantic reasoning challenges. Collectively, these design choices make TeleEgo a robust and discriminative benchmark for evaluating embodied video understanding models.
@@
Method Params Omni Streaming Memory (%) Understanding (%) Cross–Memory Reasoning (%) Overall UlM StM ET TCI LtM All II CU CmU MsR All CeR TCU CtC All Proprietary MLLMs GPT-4o [12] 47.71 47.68 34.05 39.76 34.87 42.69 78.40 60.09 53.59 47.30 60.92 33.11 57.69 58.73 45.87 48.04 Gemini-2.5-Pro [4] 46.52 49.63 30.47 35.34 39.92 42.23 71.36 60.56 49.28 47.30 57.98 28.48 50.00 52.38 40.26 46.35 Open–Source MLLMs Qwen2.5-VL [1] 8B 34.63 39.85 27.24 37.35 28.57 34.24 43.66 44.60 28.71 22.30 35.89 21.85 26.92 34.13 27.39 33.96 VideoChat-Online [11] 4B 32.26 31.78 23.66 26.51 24.37 28.91 57.28 44.13 34.93 25.68 41.76 18.54 42.06 26.92 29.04 32.46 Qwen2.5-Omni [20] 7B 26.15 29.58 21.15 26.10 20.17 25.34 32.86 32.39 23.44 17.57 27.33 15.89 23.08 24.60 20.13 25.33 MiniCPM-o [19] 8B 43.63 44.01 29.75 39.36 39.50 40.36 67.14 51.17 40.67 37.84 50.19 25.83 26.92 55.56 38.28 42.84
@@
Table 3: RTA results on the TeleEgo benchmark. Columns are grouped into three capability blocks, Memory, Understanding, and Cross–Memory Reasoning, with an All column summarizing each block and an Overall column aggregating across blocks. “Omni” denotes integrated audio–video–text perception; “Streaming” denotes native support for streaming interaction. A ✓ indicates the capability is supported, and ✗ indicates not supported. Unless otherwise specified, all systems are evaluated with synchronized video+audio+text inputs; for models without native speech recognition, we attach an ASR front end to process the audio stream.
@@
Method Params Omni Streaming Memory (min) Understanding (min) Cross–Memory Reasoning (min) Overall UlM StM ET TCI LtM All II CU CmU MsR All CeR TCU CtC All Proprietary MLLMs GPT-4o [12] 3.41 2.66 1.36 2.06 1.82 2.60 6.27 4.36 2.80 3.14 4.49 1.60 4.04 3.82 3.06 3.01 Gemini-2.5-Pro [4] 3.17 2.22 1.13 1.94 1.89 2.37 5.70 4.58 2.34 3.11 4.22 1.15 4.19 3.26 2.62 2.76 Open–Source MLLMs Qwen2.5-VL [1] 8B 1.80 1.77 1.39 1.76 1.11 1.66 2.09 2.29 1.22 0.87 1.83 0.97 1.42 1.56 1.31 1.60 VideoChat-Online [11] 4B 1.30 1.03 0.65 0.84 0.82 1.03 3.72 2.17 1.25 1.11 2.41 0.46 0.62 1.87 1.32 1.33 Qwen2.5-Omni [20] 7B 1.06 1.09 0.80 1.20 0.61 1.00 1.37 1.32 1.02 0.74 1.20 0.48 1.04 1.02 0.81 1.00 MiniCPM-o [19] 8B 2.27 1.58 1.16 1.79 1.66 1.82 4.58 3.43 2.03 2.18 3.37 0.99 1.77 3.46 2.53 2.19
Table 4: MPT results on the TeleEgo benchmark. Columns are grouped into three capability blocks, Memory, Understanding, and Cross–Memory Reasoning, with an All column summarizing each block and an Overall column aggregating across blocks. “Omni” denotes integrated audio–video–text perception; “Streaming” denotes native support for streaming interaction. A ✓ indicates the capability is supported, and ✗ indicates not supported. Unless otherwise specified, all systems are evaluated with synchronized video+audio+text inputs; for models without native speech recognition, we attach an ASR front end to process the audio stream.
@@ 4.3 Main Results
As shown in Table 3 and Table 4, under the strict streaming protocol and evidence-compliance constraints of TeleEgo, the results reveal a clear and structured pattern. Proprietary multimodal assistants (e.g., GPT-4o, Gemini-2.5-Pro) achieve strong overall performance in both RTA and MPT. However, their advantage is concentrated in the Understanding axis (e.g., GPT-4o reaches 61% in Understanding-All), while performance drops significantly in tasks requiring fine-grained temporal binding and cross-modal attribution (43% in Memory-All, 46% in Cross-Memory-All). This “semantic-strong but temporally-weak” trend is consistent across subtasks: intent inference approaches near-ceiling accuracy, whereas entity tracking and cross-entity relation inference remain the weakest, indicating that current systems heavily rely on semantic priors but struggle with timestamp alignment and instance-level grounding.
Interestingly, some open-source models with native streaming designs (e.g., MiniCPM-o) significantly close the RTA gap with proprietary systems, despite having fewer parameters. This suggests that managing temporal states and controlling output emissions may matter more than broad multimodal coverage. Conversely, models with audio-visual-text fusion but without streaming mechanisms show limited benefit in TeleEgo’s “correct-then-timed-and-verifiable” setting. This highlights that latency handling, cache scheduling, and alignment logic are the true drivers of real-time accuracy.
From a temporal persistence perspective, MPT further reveals a disconnect between what models remember and how long they remember it. Proprietary models sustain longer persistence on understanding-oriented tasks (e.g., GPT-4o achieves 6.3 minutes MPT on intent inference) but only 2–3 minutes on memory-centric tasks. Open-source models show shorter persistence across the board. This suggests that while models can compress long experiences into abstract semantic representations, they struggle to retain auditable, time-anchored evidence and dynamic entity states.
The divergence between RTA and MPT under a 5-second decision window and evidence-overlap constraints points to two complementary optimization directions: 1) Timestamp-aware temporal learning — where decoding conditioned on timestamps and calibrated silence policies improve when to respond; 2) Structured long-term memory architectures — integrating clock-indexed event keys with multimodal anchors to improve how to substantiate outputs. Overall, TeleEgo tightly couples correctness, grounding, and timing, shifting the primary bottleneck of egocentric assistants from sheer context length to verifiable alignment and real-time temporal control. This establishes a concrete and actionable frontier for future research.
@@ 5 Conclusion
In summary, we present TeleEgo, an online, omni-modal, first-person benchmark grounded in real-world use, built from continuous, multi-participant, multi-scene, multi-day recordings that align video, ambient speech/dialogue, and dual textual timelines under a unified clock, and equipped with a contract-based annotation scheme that binds each query to its required modalities and precise, time-stamped evidence spans for auditable attribution. Centered on three capability axes, Memory, Understanding, and Cross-Memory Reasoning, TeleEgo offers a fine-grained task suite and a strict streaming-only evaluation protocol: responses receive credit only if they arrive within task-specific decision windows and satisfy evidence compliance. Two complementary metrics, Real-time Accuracy, and Memory Persistence Time, jointly assess correctness, response timing, and long-horizon memory, while failure cases are decomposed into retention, retrieval, alignment, and timing to yield actionable diagnostics. We envision TeleEgo as an ecologically valid, diagnostically informative, and reproducible foundation for building first-person assistants that must remember, align, and act in real time.

### v4

@@ Abstract
Egocentric AI assistants in real-world settings must process multi-modal inputs (video, audio, text), respond in real time, and retain evolving long-term memory. However, existing benchmarks typically evaluate these abilities in isolation, lack realistic streaming scenarios, or support only short-term tasks. We introduce TeleEgo, a long-duration, streaming, omni-modal benchmark for evaluating egocentric AI assistants in realistic daily contexts. The dataset features over 14 hours per participant of synchronized egocentric video, audio, and text across four domains: work & study, lifestyle & routines, social activities, and outings & culture. All data is aligned on a unified global timeline and includes high-quality visual narrations and speech transcripts, curated through human refinement. TeleEgo defines 12 diagnostic subtasks across three core capabilities: Memory (recalling past events), Understanding (interpreting the current moment), and Cross-Memory Reasoning (linking distant events). It contains 3,291 human-verified QA items spanning multiple question formats (single-choice, binary, multi-choice, and open-ended), evaluated strictly in a streaming setting. We propose Real-Time Accuracy (RTA) to jointly capture correctness and responsiveness under tight decision windows, and Memory Persistence Time (MPT) as a forward-looking metric for long-term retention in continuous streams. In this work, we report RTA results for current models and release TeleEgo, together with an MPT evaluation framework, as a realistic and extensible benchmark for future egocentric assistants with stronger streaming memory, enabling systematic study of both real-time behavior and long-horizon memory.
@@ 1 Introduction
To address these challenges, we present TeleEgo: a long-duration, streaming, and fully multimodal benchmark grounded in real-world scenarios, purpose-built to evaluate egocentric AI assistants (see Figure 1). TeleEgo consists of synchronized video, audio, and textual data collected from multiple participants, each contributing more than 14 hours of recordings. The dataset spans four major domains: work & study, lifestyle & routines, social activities, and outings & culture. All data streams are precisely aligned to a unified global timeline and enriched with manually curated speech transcripts and visual narrations to ensure high quality and semantic clarity. Building on this foundation, TeleEgo introduces 12 diagnostic tasks covering three core ability: memory (recalling past events), understanding (interpreting the present), and cross-memory reasoning (connecting distant moments). These tasks include a total of 3,291 human-verified QA items across various formats, including single-choice, binary, multiple-choice, and open-ended questions. Each task is tied to a specific time point and a decision window, requiring models not only to answer correctly, but also to respond at the right time. All evaluations are conducted under streaming conditions, and a task is considered successful only if the model’s first correct answer falls within the allowed window. This setup prevents models from guessing too early or answering too late, offering a more accurate measure of real-time decision-making and responsiveness. We further introduce two evaluation metrics: Real-Time Accuracy, which measures whether the model produces a correct response within a given decision window, and Memory Persistence Time, a forward-looking metric that tracks how long past information remains correctly retrievable in a continuous stream without replaying evidences. In this work, we primarily report RTA on existing models and provide MPT as an evaluation protocol for future systems with stronger long-term streaming memory.
@@
M3-Bench [15] Agent Memory and Reasoning avg. 34 min Agent-centric QA Partial
@@
  • An evaluation protocol for streaming inputs, using Real-Time Accuracy metric to assess contemporary models and a forward-looking Memory Persistence Time metric to study long-term retention.
2 Related Works
Egocentric Models and Benchmarks. Egocentric (first-person) vision research has grown significantly, evolving from early single-user recordings to diverse, large-scale datasets capturing daily life from a first-person perspective. Pioneering work such as EPIC-KITCHENS [5] and its extensions like VISOR [6] introduced large-scale, object-rich video datasets in home environments. Later, Ego4D and Ego-Exo4D [7, 8] broadened the scope to include tasks like episodic memory, future prediction, and skill learning, incorporating both egocentric and exocentric views. Recent datasets have begun to explore more complex cognitive tasks and assistant-like interactions. EgoLife [21] captures long-duration recordings from a single household, while EgoThink [3] and EgoExoLearn [9] focus on segment-level reasoning and teaching-following dynamics. Other work like MM-Ego [22] and EgoTextVQA [25] targets memory and text-based understanding. However, most of these evaluations remain offline and task-specific. TeleEgo advances egocentric benchmarking by combining multi-day, multi-role, and multi-theme recordings with dual text annotations aligned to a unified global timeline. Its online evaluation protocol measures not only real-time decision-making, but also the persistence of memory over time—enabling more realistic evaluations of assistant capabilities.
@@ 2 Related Works
Omni-Modal Assistants. Recent omni-modal models seek to unify text, vision, audio, and speech for more flexible assistants. Closed-source systems such as GPT-4o [11] and Gemini 1.5/2.5 [18, 4] show strong empirical performance on speech and video, but their architectures, memory mechanisms, and streaming behavior are proprietary and largely unknown. Open-source omni-modal models have also progressed rapidly. LLaVA–OneVision [12] and InternVL-2.5 [2] support multi-image and video understanding, while Qwen2.5–Omni [20] and Baichuan–Omni [13] are optimized for audiovisual input and interactive responses. Yet they are mostly evaluated on offline, task-specific benchmarks or short clips rather than long, continuous egocentric streams. MiniCPM–o [19] moves closer to streaming by providing an online interface for token- and frame-level inputs, but relies on a growing KV cache without effective long-term memory control. Beyond single-model interfaces, M3-Agent [15] introduces a multimodal agent with explicit long-term memory, organizing episodic and semantic information in an entity-centric, multimodal form. However, it still assumes segment-wise video input instead of second-by-second egocentric streams. Despite these advances, few works test whether omni-modal assistants can both respond accurately under time-sensitive decision windows and retain information over long temporal spans in realistic, continuous settings. TeleEgo combines egocentric recordings, streaming video contexts, and real-time memory-oriented evaluation to provide a unified testbed bridging egocentric datasets, streaming benchmarks, and multimodal models.
@@ 3.4 Benchmark QA Annotation
As shown in Figure 4, the TeleEgo benchmark is hierarchically structured along three key cognitive dimensions—memory, understanding, and reasoning—with twelve subcategories providing finer granularity. The final dataset contains 3,291 verified QA instances across all four formats (see Table 2). Subcategory distributions demonstrate balanced coverage across temporal, causal, and semantic reasoning challenges. Collectively, these design choices make TeleEgo a robust and discriminative benchmark for evaluating video understanding models.
@@
Method Params Omni Streaming Memory (%) Understanding (%) Cross–Memory Reasoning (%) Overall UlM StM ET TCI LtM All II CU CmU MsR All CeR TCU CtC All GPT-4o [11] 42.31 40.58 31.91 47.37 52.78 42.18 81.81 81.58 45.71 50.00 66.67 44.44 40.00 45.00 44.23 48.94 Gemini-2.5-Pro [4] 49.04 45.59 34.04 47.37 44.44 45.05 63.64 55.26 37.14 28.57 48.03 40.74 40.00 45.00 42.31 45.55
@@
Table 3: RTA results of proprietary MLLMs on TeleEgo (single-video evaluation). Columns are grouped into three capability blocks: Memory, Understanding, and Cross–Memory Reasoning, with an All column summarizing each block and an Overall column aggregating across blocks. “Omni” denotes integrated audio–video–text perception; “Streaming” denotes native support for streaming interaction. A ✓ indicates the capability is supported, ✗ indicates not supported, and “–” indicates unknown. Due to the extensive video duration ($\sim$14 hours) and API latency, both models were evaluated on the same participant’s video for fair comparison.
@@
Method Params Omni Streaming Memory (%) Understanding (%) Cross–Memory Reasoning (%) Overall UlM StM ET TCI LtM All II CU CmU MsR All CeR TCU CtC All VideoChat-Online [10] 4B 30.74 26.29 19.35 29.84 22.78 26.69 57.35 44.13 35.44 27.89 42.34 18.79 32.00 41.60 29.43 31.28 Qwen2.5-VL [1] 8B 45.42 42.01 31.18 37.90 33.33 39.66 66.35 53.99 47.57 41.50 53.28 32.89 48.00 52.00 42.14 43.67 Qwen2.5-Omni [20] 7B 44.21 42.75 35.48 41.13 37.97 41.20 72.51 61.50 50.00 48.98 59.07 37.58 56.00 61.60 49.16 46.96
Table 4: RTA results of pseudo-streaming open-source MLLMs on TeleEgo. These models lack cross-call memory mechanisms and process each input unit independently. They are categorized as “pseudo-streaming” because they lack native support for streaming interaction. To test the models’ intrinsic memory capabilities, they are evaluated under a strict streaming protocol where only the current frame and audio are provided at each second.
Memory (%) Understanding (%) Cross–Memory Reasoning (%) Method Params Omni Streaming Session Length UlM StM ET TCI LtM All II CU CmU MsR All CeR TCU CtC All Overall 1sec 51.99 51.60 35.13 47.58 44.30 47.54 75.83 61.50 51.94 44.22 59.59 28.19 48.00 64.80 45.15 50.60 1min 50.60 55.53 37.28 53.63 51.05 50.11 80.57 69.48 56.80 51.02 65.64 33.56 36.00 66.40 47.49 54.10 3min 48.19 52.58 37.28 49.19 45.99 47.31 73.93 63.38 53.88 40.14 59.33 32.89 32.00 66.40 46.82 50.57 5min 43.00 44.23 33.69 47.58 37.55 41.71 63.51 56.34 44.17 43.54 52.64 23.49 32.00 56.80 38.13 44.34 MiniCPM-o [19] 8B 10min 28.50 28.01 23.66 30.65 26.16 27.60 41.71 40.38 34.95 29.93 37.32 15.44 36.00 37.60 26.42 30.15
Table 5: RTA results of MiniCPM-o with varying session lengths on TeleEgo. Session length determines how long the model retains memory of previously observed frames before reset. “1s” corresponds to no memory (each call is independent). Longer sessions allow more historical context but may cause attention dilution and generation quality degradation.
@@ 4.3 Main Results
We group our RTA analysis into three model categories, each under a distinct evaluation setting.
Proprietary API Models.
GPT-4o and Gemini-2.5-Pro are evaluated via API calls, with their internal implementations opaque. Due to the long video duration ($\sim$14 hours per participant) and API latency, we report RTA results on a single participant (Table 3). GPT-4o shows strong Understanding performance (66.67%), especially on Intent Inference (81.81%) and Causal Understanding (81.58%). This strength aligns with the nature of Understanding tasks, which mainly require semantic reasoning and commonsense knowledge rather than precise temporal localization, allowing the model to exploit its general-purpose priors. In contrast, performance drops on Memory (42.18%) and Cross-Memory Reasoning (44.23%), where fine-grained temporal binding and evidence grounding are crucial.
Pseudo-Streaming Open-Source Models.
VideoChat-Online, Qwen2.5-VL, and Qwen2.5-Omni lack genuine cross-call memory mechanisms (Table 4). VideoChat-Online’s Pyramid Memory Bank exists only within a single inference call and is reset at the end of each call. Qwen2.5-VL and Qwen2.5-Omni support multi-turn dialogue by naively accumulating raw conversation history, but provide no explicit streaming interface or memory management. In our evaluation, we deliberately avoid feeding uncompressed historical context into the models, since this merely shifts the memory burden to the context window rather than testing intrinsic memory capability. Instead, we adopt a strict streaming protocol: at each timestep, the model receives only the current frame and audio. Under this setting, Understanding tasks remain manageable, while Memory and Cross-Memory Reasoning become much harder. Among the three, Qwen2.5-Omni achieves the highest overall RTA (46.96%), primarily because it has stronger base model and avoids the ASR transcription errors.
Streaming Model: MiniCPM-o.
MiniCPM-o is the only model in our study with an explicit streaming interface that accepts chunk-wise incremental input (Table 5). However, it lacks dynamic memory management: the KV cache grows without compression, eviction, or sliding window. This causes two issues in practice: GPU memory overflow as video length increases, and attention dilution that degrades generation quality in long sessions. We therefore reset the session at fixed intervals. RTA results indicate an optimal session length of about 1 minute (54.10% overall), with poorer performance at both shorter (50.60% at 1s) and longer intervals (30.15% at 10min). At longer intervals, the model often produces incoherent or off-topic responses, confirming that naive KV cache accumulation without memory management is insufficient for sustained streaming. This highlights an open challenge: efficiently managing long-term memory under bounded computational resources remains unresolved for streaming MLLMs.
4.4 MPT in Current Models
We proposed Memory Persistence Time (MPT) to measure how long a model retains information after initially answering correctly. However, the current evaluation does not include MPT results, as none of the tested models are suitable for this metric. For pseudo-streaming models (VideoChat-Online, Qwen2.5-VL, Qwen2.5-Omni), each inference call is independent with no cross-call memory. In this setting, MPT is ill-defined: if a memoryless model answers correctly in both the initial and recall phases, we cannot tell whether it “remembered” the information or simply “guessed correctly twice”. For MiniCPM-o, although the KV cache can persist across calls, the lack of dynamic memory management forces us to reset sessions every few minutes. When an initial test and its recall test span a session reset, the model’s memory is cleared in between. Any recall failure may then result from genuine forgetting or from the reset itself, making MPT measurements unreliable.
MPT is therefore a forward-looking metric that aligns with TeleEgo’s goal of evaluating realistic long-term egocentric assistants. At present, models with true long-horizon streaming memory are still rare: few can run continuously for hours without resets while maintaining stable generation quality. As stronger streaming MLLMs emerge, MPT will become central to measuring how long information is retained beyond the moment of first use. TeleEgo, together with our released MPT evaluation framework and code, provides a unified setting to assess both real-time correctness and long-term memory persistence, paving the way for future work on practical egocentric AI assistants.
@@ 5 Conclusion
We present TeleEgo, an online, omni-modal, first-person benchmark grounded in real-world use. It is built from continuous, multi-participant, multi-day recordings that align egocentric video, ambient speech/dialogue, and textual timelines under a unified global clock. A contract-based annotation scheme binds each query to its required modalities and to precise, time-stamped evidence spans, enabling auditable attribution. Centered on three capability axes—Memory, Understanding, and Cross-Memory Reasoning—TeleEgo employs a strict streaming-only evaluation protocol where models consume inputs once and responses must arrive within task-specific decision windows. We employ Real-Time Accuracy as the main metric in our experiments, capturing both correctness and timing under online decision constraints. In addition, we introduce Memory Persistence Time as a forward-looking protocol for probing long-horizon retention in continuous streams, and release an accompanying evaluation framework for future systems with stronger long-term streaming memory, along with failure breakdowns over retention, retrieval, alignment, and timing. We envision TeleEgo as an ecologically valid foundation for developing first-person assistants that must remember, align, and act in real time.